# Your Turn: Real-World Turning Angle Estimation for Parkinson’s Disease Severity Assessment

### v1

@@ Abstract
People with Parkinson’s Disease (PD) often experience progressively worsening gait, including changes in how they turn around, as the disease progresses. Existing clinical rating tools are not capable of capturing hour-by-hour variations of PD symptoms, as they are confined to brief assessments within clinic settings, leaving gait performance outside these controlled environments unaccounted for. Measuring real-world gait turning angles continuously and passively is a component step towards using gait characteristics as sensitive indicators of disease progression in PD. This paper presents a deep learning-based approach to automatically quantify turning angles by extracting 3D skeletons from videos and calculating the rotation of hip and knee joints. We utilise state-of-the-art human pose estimation models, Fastpose and Strided Transformer, on a total of 1386 turning video clips from 24 subjects (12 people with PD and 12 healthy control volunteers), trimmed from a PD dataset of unscripted free-living videos in a home-like setting (Turn-REMAP). We also curate a turning video dataset, Turn-H3.6M, from the public Human3.6M human pose benchmark with 3D ground truth, to further validate our method. Previous gait research has primarily taken place in clinics or laboratories evaluating scripted gait outcomes, but this work focuses on real-world settings where complexities exist, such as baggy clothing and poor lighting. Due to difficulties in obtaining accurate ground truth data in a free-living setting, we quantise the angle into the nearest bin $45^{\circ}$ based on the manual labelling of expert clinicians. Our method achieves a turning calculation accuracy of 41.6%, a Mean Absolute Error (MAE) of 34.7°, and a weighted precision (WPrec) of 68.3% for Turn-REMAP. On Turn-H3.6M, it achieves an accuracy of 73.5%, an MAE of 18.5°, and a WPrec of 86.2%. This is the first work to explore the use of single monocular camera data to quantify turns by PD patients in a home setting. All data and models are publicly available, providing a baseline for turning parameter measurement to promote future PD gait research.
@@ 1 Introduction
Parkinson’s disease (PD) is a progressive neurodegenerative movement disorder, characterised by symptoms such as slowness of movement and gait dysfunction [Jankovic2008] which fluctuate across the day but progress slowly over the years [Holden2018]. Currently, treatment of PD relies on therapies which improve symptoms. There are no treatments available which modify the course of the underlying disease (so-called disease-modifying treatments, or DMTs), despite there being multiple putative DMTs showing promise in laboratory studies [Lang2013]. One reason for the slow development of DMTs is the dearth of sensitive, frequent, objective biomarkers to enhance the current gold-standard clinical rating scale [Goetz2008] to measure the progression of PD. This gold-standard clinical rating scale, the Movement Disorders Society-sponsored revision of the Unified Parkinson’s Disease Rating Scale (MDS-UPDRS) [Goetz2008], includes subjective questionnaires concerning gait and mobility experiences, along with clinicians’ ratings of scripted activities performed by the participants. The assessments typically occur within clinical settings over short durations, offering only a ”snapshot” of symptoms which vary on an hourly basis. It also has limitations including its non-linear and discontinuous scoring system, the inter-rater variability [Post2005] and Hawthorne effect [Paradis2017] of being observed on how someone mobilises [Robles-Garcia2015, Morberg2018]. Gait and turning abnormalities are common features of PD with over half of the patients reporting difficulties with turning [Bloem2001] – when someone moves round on their axis while upright, changing the direction they face. Turning changes associated with PD include the ’en bloc’ phenomenon where upper and lower body segments turn simultaneously [Spildooren2013], a longer duration of turn, less accurate turn completion, a narrower base of support [Mellone2016] and the use of ’step turns’ rather than ’pivot turns’ [Hulbert2015]. More than 40% of daily steps are during turning [Glaister2007] and turning abnormalities can predispose to falls, thus turning parameters could potentially be used as measures predicting the time to falls in a patient with PD [Bloem2001]. Furthermore, if a fall happens during turning, it is up to 8 times more likely to result in a hip fracture [Cummings1994]. In unmedicated early-stage PD, gait parameters from turning are more sensitive to change compared to straight-ahead gait outcomes [Zampieri2010], making measuring aspects of turns potentially of specific use in clinical trials of disease-modifying interventions which typically recruit recently diagnosed patients [Stephenson2021]. People with PD turn differently when being watched by a clinician [Morgan2022PRD], so measuring turning passively in a real-world setting (Figure 1) could give information about mobility not captured by face-to-face assessments in the clinic.
@@
Figure 1: Examples of real-world turning activity in REMAP [Morgan2023_REMAPOpen]. – Sample frames (first row) in the living room and (second row) in a kitchen.
@@
    Utilising human pose estimation models, we propose a pipeline to estimate turning angles from single-view RGB videos and validate this pipeline on our proposed datasets. This is the first work to estimate turning angles from real-world free-living video data captured from people living with PD.
@@ 2 Literature review
Another inherent limitation of wearable-based methods is that they can only provide kinematics parameters on a few body parts, rather than a holistic view of the position and orientation of the entire body. Furthermore, it is shown in [Rehman2020] that placing wearables on different locations of the body (head, neck, lower back and ankle) causes inconsistency in estimated turning angles. Video-based markerless approaches [kidzinski2020deep, shin2021quantitative, stenum2021two] present a passive and less obtrusive solution to these innate problems of wearable-based approaches. However, compared to marker-based approaches, the accuracy [wade2022applications] of joint angle estimation is not yet good enough for clinical application. The work in [shah2020inertial, wade2022applications] has shown that the reported performances are inconsistent and hard to reproduce outside laboratory environments and among different patient cohorts, as they often use off-the-shelf human pose analysis software and hardware on experiments set up in restricted laboratories with scripted activities. To develop and validate robust video-based gait analysis algorithms, the challenge lies in acquiring videos that are representative enough across different patients in different scenarios including clinics, hospitals and homes. We gathered and annotated a dedicated real-world video dataset for turning angle estimation to complement the existing gait analysis research for PD.
@@
We argue that for PD, it is important to accurately measure absolute gait parameters like turning angle, in a real-world, non-hospital setting, to build sufficiently sensitive markers to reliably track changes in gait throughout the day and over the years. To this end, most of the previous pose-based research on gait analysis with angular measurement focuses on sagittal [kidzinski2020deep, abe2021openpose] or coronal [shin2021quantitative, tang20222d] joint angles. Only [cao2019human, kondragunta2019estimation, shin2021quantitative] use pose estimation algorithms on turning analysis, but all three are solely for turn detection, or other gait metrics like step length rather than turning angle. Also, other sensors like 3D depth cameras [clark2019three] have been used for human pose estimation using depth maps or point clouds. While such an approach is optimal in some applications like virtual reality, there are many unresolved challenges such as being unable to handle self-occlusions or multi-person detection [xu2021review]. In the complex, unscripted scenario of monitoring everyday activities, it may not be suitable to use depth sensors alone; however, combining it with RGB data could potentially lead to better results.
@@ 7 Discussion
Although the overall measurement accuracy of our Turn-REMAP dataset is not yet robust enough for clinical diagnosis, it establishes a baseline for future, passive, video-based analysis of turning movements in indoor, free-living environments. Our manual examination of incorrectly classified video clips and their corresponding 3D skeletons revealed that depth reconstruction ambiguity is a significant factor [martinez2017simple, zhang2022mixste] that significantly affects the accurate calculation of turning angles. Recovering the missing depth from a 2D image is inherently an ill-posed problem as infinitely many 3D poses can project to the same 2D skeleton. Despite our utilised models being pretrained on large-scale laboratory 3D motion data, generalising this performance to reconstruct unseen poses in our real-world PD turning dataset remains a substantial challenge.
@@
The statistics reveal challenges in generalising our pretrained human pose estimation model from the lab-based Turn-H3.6M dataset to the diverse, real-world Turn-REMAP dataset. Different global position distributions [chai2023global], camera parameters [zhan2022ray3d], and diverse human body sizes and shapes, as well as articulated movements [gong2021poseaug, gholami2022adaptpose] highlight the need to enhance model robustness and adaptability to better handle real-world variability. To bridge this gap and enhance the ability to generalise to new, unseen data, it is crucial to implement domain adaptation strategies in deep learning models and conduct cross-dataset validation.
@@ 8 Conclusion and Future Work
Automatically computing turning angles in a free-living environment is foundational for future longitudinal, in-home monitoring of PD. There are many potential avenues to build upon our work for more accurate turning angle estimation. Although Turn-REMAP and Turn-H3.6M only consist of trimmed turning clips, our methods can be extended to untrimmed videos. We could also infer additional turning metrics such as turning speed from the estimated turning angle. These metrics can be used to classify PD and control subjects, infer clinical rating scores of disease severity, or assess on/off medication status in free-living video recordings. Another extension for more accurate turning angle computation could be to replace our skeleton model with other models, such as via HMR (Human Mesh Recovery) [goel2023humans] which could offer additional parameters for turning angle estimation.
@@ Acknowledgments
The authors gratefully thank the study participants for their time and effort in participating in this research. We also acknowledge the local Parkinson’s and other Movement Disorders Health Integration Team (Patient and Public Involvement Group) for their assistance at each step of the study design. This work was supported by the SPHERE Next Steps Project funded by the EPSRC (grant EP/R005273/1),the Elizabeth Blackwell Institute for Health Research at the University of Bristol, the Wellcome Trust Institutional Strategic Support Fund (grant 204813/Z/16/Z), Cure Parkinson’s Trust (grant AW021), and by IXICO (grant R101507-101). Dr Jonathan de Pass and Mrs Georgina de Pass also made a charitable donation to the University of Bristol through the Development and Alumni Relations Office to support research into Parkinson’s Disease. The first author was supported by the Engineering and Physical Sciences Research Council (EPSRC) Digital Health and Care Centre for Doctoral Training (CDT) at the University of Bristol (UKRI Grant No. EP/S023704/1).

### v2

# Your Turn: At Home Turning Angle Estimation for Parkinson’s Disease Severity Assessment

@@ Abstract
People with Parkinson’s Disease (PD) often experience progressively worsening gait, including changes in how they turn around, as the disease progresses. Existing clinical rating tools are not capable of capturing hour-by-hour variations of PD symptoms, as they are confined to brief assessments within clinic settings, leaving gait performance outside these controlled environments unaccounted for. Measuring turning angles continuously and passively is a component step towards using gait characteristics as sensitive indicators of disease progression in PD. This paper presents a deep learning-based approach to automatically quantify turning angles by extracting 3D skeletons from videos and calculating the rotation of hip and knee joints. We utilise state-of-the-art human pose estimation models, Fastpose and Strided Transformer, on a total of 1386 turning video clips from 24 subjects (12 people with PD and 12 healthy control volunteers), trimmed from a PD dataset of unscripted free-living videos in a home-like setting (Turn-REMAP). We also curate a turning video dataset, Turn-H3.6M, from the public Human3.6M human pose benchmark with 3D ground truth, to further validate our method. Previous gait research has primarily taken place in clinics or laboratories evaluating scripted gait outcomes, but this work focuses on free-living home settings where complexities exist, such as baggy clothing and poor lighting. Due to difficulties in obtaining accurate ground truth data in a free-living setting, we quantise the angle into the nearest bin $45^{\circ}$ based on the manual labelling of expert clinicians. Our method achieves a turning calculation accuracy of 41.6%, a Mean Absolute Error (MAE) of 34.7°, and a weighted precision (WPrec) of 68.3% for Turn-REMAP. On Turn-H3.6M, it achieves an accuracy of 73.5%, an MAE of 18.5°, and a WPrec of 86.2%. This is the first work to explore the use of single monocular camera data to quantify turns by PD patients in a home setting. All data and models are publicly available, providing a baseline for turning parameter measurement to promote future PD gait research.
@@ 1 Introduction
Parkinson’s disease (PD) is a progressive neurodegenerative movement disorder, characterised by symptoms such as slowness of movement and gait dysfunction [Jankovic2008] which fluctuate across the day but progress slowly over the years [Holden2018]. Currently, treatment of PD relies on therapies which improve symptoms. There are no treatments available which modify the course of the underlying disease (so-called disease-modifying treatments, or DMTs), despite there being multiple putative DMTs showing promise in laboratory studies [Lang2013]. One reason for the slow development of DMTs is the dearth of sensitive, frequent, objective biomarkers to enhance the current gold-standard clinical rating scale [Goetz2008] to measure the progression of PD. This gold-standard clinical rating scale, the Movement Disorders Society-sponsored revision of the Unified Parkinson’s Disease Rating Scale (MDS-UPDRS) [Goetz2008], includes subjective questionnaires concerning gait and mobility experiences, along with clinicians’ ratings of scripted activities performed by the participants. The assessments typically occur within clinical settings over short durations, offering only a ”snapshot” of symptoms which vary on an hourly basis. It also has limitations including its non-linear and discontinuous scoring system, the inter-rater variability [Post2005] and Hawthorne effect [Paradis2017] of being observed on how someone mobilises [Robles-Garcia2015, Morberg2018]. Gait and turning abnormalities are common features of PD with over half of the patients reporting difficulties with turning [Bloem2001] – when someone moves round on their axis while upright, changing the direction they face. Turning changes associated with PD include the ’en bloc’ phenomenon where upper and lower body segments turn simultaneously [Spildooren2013], a longer duration of turn, less accurate turn completion, a narrower base of support [Mellone2016] and the use of ’step turns’ rather than ’pivot turns’ [Hulbert2015]. More than 40% of daily steps are during turning [Glaister2007] and turning abnormalities can predispose to falls, thus turning parameters could potentially be used as measures predicting the time to falls in a patient with PD [Bloem2001]. Furthermore, if a fall happens during turning, it is up to 8 times more likely to result in a hip fracture [Cummings1994]. In unmedicated early-stage PD, gait parameters from turning are more sensitive to change compared to straight-ahead gait outcomes [Zampieri2010], making measuring aspects of turns potentially of specific use in clinical trials of disease-modifying interventions which typically recruit recently diagnosed patients [Stephenson2021]. People with PD turn differently when being watched by a clinician [Morgan2022PRD], so measuring turning passively in uncontrolled home settings (Figure 1) could give information about mobility not captured by face-to-face assessments in the clinic.
@@
Figure 1: Examples of free-living home scenario turning activity in REMAP [Morgan2023_REMAPOpen]. Sample frames (first row) in the living room and (second row) in a kitchen.
@@
    Utilising human pose estimation models, we propose a pipeline to estimate turning angles from single-view RGB videos and validate this pipeline on our proposed datasets. This is the first work to estimate turning angles from natural free-living video data captured from people living with PD.
@@ 2 Literature review
Another inherent limitation of wearable-based methods is that they can only provide kinematics parameters on a few body parts, rather than a holistic view of the position and orientation of the entire body. Furthermore, it is shown in [Rehman2020] that placing wearables on different locations of the body (head, neck, lower back and ankle) causes inconsistency in estimated turning angles. Video-based markerless approaches [kidzinski2020deep, shin2021quantitative, stenum2021two] present a passive and less obtrusive solution to these innate problems of wearable-based approaches. However, compared to marker-based approaches, the accuracy [wade2022applications] of joint angle estimation is not yet good enough for clinical application. The work in [shah2020inertial, wade2022applications] has shown that the reported performances are inconsistent and hard to reproduce outside laboratory environments and among different patient cohorts, as they often use off-the-shelf human pose analysis software and hardware on experiments set up in restricted laboratories with scripted activities. To develop and validate robust video-based gait analysis algorithms, the challenge lies in acquiring videos that are representative enough across different patients in different scenarios including clinics, hospitals and homes. We gather and annotate a dedicated, free-living video data set to estimate turning angles to complement existing research on gait analysis for PD.
@@
We argue that for PD, it is important to accurately measure absolute gait parameters like turning angle, in a natural, non-hospital setting, to build sufficiently sensitive markers to reliably track changes in gait throughout the day and over the years. To this end, most of the previous pose-based research on gait analysis with angular measurement focuses on sagittal [kidzinski2020deep, abe2021openpose] or coronal [shin2021quantitative, tang20222d] joint angles. Only [cao2019human, kondragunta2019estimation, shin2021quantitative] use pose estimation algorithms on turning analysis, but all three are solely for turn detection, or other gait metrics like step length rather than turning angle. Also, other sensors like 3D depth cameras [clark2019three] have been used for human pose estimation using depth maps or point clouds. While such an approach is optimal in some applications like virtual reality, there are many unresolved challenges such as being unable to handle self-occlusions or multi-person detection [xu2021review]. In the complex, unscripted scenario of monitoring everyday activities, it may not be suitable to use depth sensors alone; however, combining it with RGB data could potentially lead to better results.
@@ 7 Discussion
Although the overall measurement accuracy of our Turn-REMAP dataset is not yet robust enough for clinical diagnosis, it establishes a baseline for future, passive, video-based analysis of turning movements in indoor, free-living environments. Our manual examination of incorrectly classified video clips and their corresponding 3D skeletons revealed that depth reconstruction ambiguity is a significant factor [martinez2017simple, zhang2022mixste] that significantly affects the accurate calculation of turning angles. Recovering the missing depth from a 2D image is inherently an ill-posed problem as infinitely many 3D poses can project to the same 2D skeleton. Despite our utilised models being pretrained on large-scale laboratory 3D motion data, generalising this performance to reconstruct unseen poses in our in-the-wild PD turning dataset remains a substantial challenge.
@@
The statistics reveal challenges in generalising our pretrained human pose estimation model from the lab-based Turn-H3.6M dataset to the diverse, in-the-wild Turn-REMAP dataset. Different global position distributions [chai2023global], camera parameters [zhan2022ray3d], and diverse human body sizes and shapes, as well as articulated movements [gong2021poseaug, gholami2022adaptpose] highlight the need to enhance model robustness and adaptability to better handle real-world variability. To bridge this gap and enhance the ability to generalise to new, unseen data, it is crucial to implement domain adaptation strategies in deep learning models and conduct cross-dataset validation.
@@ 8 Conclusion and Future Work
Automatically computing turning angles in a free-living environment is foundational for future longitudinal, in-home monitoring of PD. There are many potential avenues to build upon our work for more accurate turning angle estimation. Although Turn-REMAP and Turn-H3.6M only consist of trimmed turning clips, our methods can be extended to untrimmed videos. We could also infer additional turning metrics such as turning speed from the estimated turning angle. These metrics can be used to classify PD and control subjects, infer clinical rating scores of disease severity, or assess on/off medication status in free-living video recordings. Another extension for more accurate turning angle computation could be to replace our skeleton model with other models, such as via Human Mesh Recovery [goel2023humans] which could offer additional parameters for turning angle estimation.
@@ Acknowledgments
The authors gratefully thank the study participants for their time and effort in participating in this research. We also acknowledge the local Parkinson’s and other Movement Disorders Health Integration Team (Patient and Public Involvement Group) for their assistance at each step of the study design. This work was supported by the SPHERE Next Steps Project funded by the EPSRC (grant EP/R005273/1),the Elizabeth Blackwell Institute for Health Research at the University of Bristol, the Wellcome Trust Institutional Strategic Support Fund (grant 204813/Z/16/Z), Cure Parkinson’s Trust (grant AW021), and by IXICO (grant R101507-101). Dr Jonathan de Pass and Mrs Georgina de Pass also made a charitable donation to the University of Bristol through the Development and Alumni Relations Office to support research into Parkinson’s Disease. The first author was supported by the Engineering and Physical Sciences Research Council Digital Health and Care Centre for Doctoral Training at the University of Bristol (UKRI Grant No. EP/S023704/1).